OpenStudy (anonymous):

5x^2=-45x? Solve for x.

OpenStudy (theviper):

\[5x^2=-45x\] \[=>5x^2+45x=0\]Take \(5x\) as common \[=>5x(x+9)=0\]\[=>x+9=0\]\[=>x=?\]

OpenStudy (anonymous):

Thank you. So x is either 0 or -9?

OpenStudy (anonymous):

yes it's both

OpenStudy (anonymous):

Okay, thanks!